(1+1/tanA)(1+1/cotA)=1/sinA-sinA
Prove that following
 

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
 (1 + tan A/ tanA) (1 + cot A/ cotA)
 
 = 1 + tan A*cot A + tanA + cotA
 
 = 2 + sinA / cosA + cosA/sinA
 
 = 2 + 1 / sinA cos A
 
 = 2 + 2 / sin2A
 
Regards
Arun